• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 20:40:21 on October 22, 2010

    Entry number 336203

    keyword=Status and Runplan for remainder of swing and owl shift

    After some issues returning beam to the hall after our Controlled Access,
    we are back in business. We have just started our first run for the low
    energy Kin-3 point.
    The runplan for tonight (swing and owl) will be to continue taking
    production data at these kinematics.
    Target 15 cm LH2
    CODA configuration: DVCSTopBot, ps5=1, all others=0 
    Beam current = 2 uA (this givs us about the same dead time that we had at
    1 uA at the last setting).
    At the end of each run, please do a full replay, log the ROC17 and ROC14
    DAQ windows, and log the scaler and CODA rates.
    Longer term items to be addressed:
    1. Moller measurement at 4 pass: the timing and need for this will depend
    on Hall C developments.
    2. ARC energy measurement: this may happen Wednesday.
    3. Compton tune: the present status is unclear. If the Compton rates are
    too bad, we may need some dedicated tuning.